OpenStudy (gabylovesyou):

Match each percent problem with its solution. 3 is what percent of 12? 12 is 4% of what number? 4% of what number is 3? What is 30% of 120? A. 36 B. 25 C. 300 D. 75

OpenStudy (gabylovesyou):

@brinethery @bobbylynn @badi @BillRosie Please help :D @amorfide

OpenStudy (anonymous):

30% of 120=36 4% of 75 is 3 12 is 4% of 300 3 is 25% of 12

OpenStudy (amorfide):

3/12=1/4=0.25 x 100 = 25% 12/4=3 x 100 = 300 3x100= 300 / 4 =75 120/10=12 x 3=36

OpenStudy (anonymous):

3 is 25 percent of 12 and 12 is 4 percent of 300 and 3 is 4 percent of 75 and 36 is 30 percent of 120

OpenStudy (gabylovesyou):

Thank you!! :D

OpenStudy (anonymous):

Welcome ((if you are indeed talking to me...)) XD

OpenStudy (gabylovesyou):

lol im talking to all of you ;)

OpenStudy (anonymous):

Oh haha okay XD